OpenStudy (anonymous):

The probability of winning a game is 15%. If you play 20 times, how many times should you expect to win?

OpenStudy (anonymous):

multiply the number of games played with probability of winning a game. so here it is 20*15/100 games= 3 games always remember, if n is the number of trials and p is the probability of an event occurring then expected number of occurrence=np

OpenStudy (imstuck):

I set it up as a ratio. If you win 15% of the time, that means that you will win 15 out every 100 times you play (because percentages are based on 100). So, if you win 15 out of 100, how many will you win out of 20?\[\frac{ 15 }{ 100 }=\frac{ x }{ 20 }\]The number of wins is on top and the number of times played is on bottom. You can do cross multiplication to solve this for x.

OpenStudy (anonymous):

Wow! Thank you guys for the great responses. I now know two new ways to figure out these problems. I can never figure them out.